# Rubric Is All You Need: Enhancing LLM-based Code Evaluation With Question-Specific Rubrics

### v1

@@ 7. Limitations and Future Work
Although this study provides valuable insights into rubric-based code evaluation, certain limitations must be acknowledged which also advocate avenues for future work. Firstly, we didn’t explore variability among various LLMs and mainly worked with GPT-4o. Different training methodologies may influence effectiveness and accuracy of code evaluation and a future research could compare multiple models. We also didn’t explore the differences between open-sources and closed-source LLMs. Our analysis is limited to programming questions in Java and do not test other widely-used languages like Python, C++. We focussed on 2 intermediary courses which had single file code assignments. A future research could delve into advanced courses which require multiple files handling. The rubrics which can be used for evaluation, can be of various granularities ranging from low to medium to fine. A detailed study on their effectiveness and comparison is another future avenue for research.
@@ 8. Conclusion
In this paper, we explored how effectively can an LLM-based grader evaluate student code using question-specific rubric compared to question-agonistic rubric. To facilitate our evaluation, we introduced two novel datasets focused on DSA and OOP, encompassing solutions of varying correctness levels. Using these datasets, we assessed our proposed techniques against existing approaches using exisitng and a new evaluation metric, Leniency. Our findings highlight the limitations of existing evaluation techniques and demonstrate the potential benefits of question-specific rubrics in enhancing logical assessment of code. This work not only provides new insights into LLM-based code evaluation but also lays the groundwork for future research into refining evaluation methods and expanding dataset availability to improve automated code assessment.
@@ 9. Acknowledgments
The authors wish to acknowledge the use of ChatGPT/Claude in the writing of this paper. This tool was used to improve the structure, presentation and grammar of the paper. The paper remains an accurate representation of the authors’ underlying work and novel intellectual contributions.

### v2

@@ 8. Limitations and Future Work
Although this study provides valuable insights into rubric-based code evaluation, certain limitations must be acknowledged which also advocate avenues for future work. Firstly, we didn’t explore variability among various LLMs and mainly worked with GPT-4o. Different training methodologies may influence effectiveness and accuracy of code evaluation and a future research could compare multiple models. We also didn’t explore the differences between open-sources and closed-source LLMs. Our analysis is limited to programming questions in Java and do not test other widely-used languages like Python, C++. We focussed on two intermediary courses which had single file code assignments. A future research could delve into advanced courses which require multiple files handling. The rubrics can be of various granularities ranging from low to medium to fine. A detailed study on their effectiveness and comparison is another future avenue for research.
@@ 9. Conclusion
In this paper, we explored how effectively can an LLM-based grader evaluate student code using a question-specific rubric compared to a question-agnostic rubric. To facilitate our evaluation, we introduced two novel datasets focused on DSA and OOP, encompassing solutions of varying correctness levels. Using these datasets, we assessed our proposed techniques against existing approaches using exisitng and a new evaluation metric, Leniency. Our findings highlight the limitations of existing evaluation techniques and demonstrate the potential benefits of question-specific rubrics in enhancing logical assessment of code. This work not only provides new insights into LLM-based code evaluation but also lays the groundwork for future research into refining evaluation methods and expanding dataset availability to improve automated code assessment.
@@ 10. Acknowledgments
This research was carried out in part with support from the New Faculty Seed Grant, Birla Institute of Technology and Science (BITS), Pilani (Grant Ref. N4/24/1004).
The authors also wish to acknowledge the use of ChatGPT/Claude in the writing of this paper. This tool was used to generate ideas regarding the presentation of tables and figures in the paper, and to improve the written grammar. The paper remains an accurate representation of the authors’ underlying work and novel intellectual contributions.